09.00 10.00 11.00 12.00 13.00 14.00 15.00 16.00 17.00 18.00 09.00 10.00 11.00 12.00 13.00 14.00 15.00 16.00 17.00 18.00 Monday Lecture 1: Lecture 2: Coffee and tea Lecture 3:

Lunch

Lecture 4: Lecture 5: Coffee and tea Lecture 6:

Tuesday Lecture 7: Lecture 8: Coffee and tea Lecture 9:

Lunch

Lecture 10: Lecture 11: Coffee and tea Lecture 12:

Wednesday Lecture 13: Lecture 14: Coffee and tea Lecture 15: Lunch

Discussion Thursday Lecture 16: Lecture 17: Coffee and tea Lecture 18:

Lunch

Lecture 19: Lecture 20: Coffee and tea Lecture 21:

Friday Avramov, I Green, I Coffee and tea Huisgen-Zimmermann, I

Lunch

Keller, I Happel, I Coffee and tea Happel, II Saturday Avramov, II Green, II Coffee and tea Huisgen-Zimmermann, II

Lunch

Keller, II Happel, III Coffee and tea Ringel, I

Sunday Avramov, III Keller, III Coffee and tea Ringel, II Lunch